OpenStudy (anonymous):

Someone help me out? I need to solve for X. 2x²+14x-60=0

OpenStudy (campbell_st):

start by dividing every term by 2 then you get \[x^2 + 7x - 30 = 0\] this can be factorised.... find the factors of -30 that add to 7 the larger factor is positive and the smaller is negative

OpenStudy (anonymous):

-3 and 10?

OpenStudy (campbell_st):

they are the factors so you have (x -3)(x+10) = 0 for the solutions you now need to solve x - 3 = 0 and x + 10 = 0

OpenStudy (anonymous):

x=3 and x=-10? I'm really confused right now lol.

OpenStudy (campbell_st):

ok... so factoring allows you rewrite the quadratic equation in a simpler form, the product of 2 binomials. (x -3)(x+10) = 0 if either binomial is equal to zero, the whole equation will be zero... and true. so thats why you let each binomial equal zero then solve it x - 3= 0..... means x = 3 is a solution to the quadratic. x + 10 - 0 means x = -10 is a solution to the quadratic you can substitute into the original equation to check 2(3)^2 + 14(3) - 60 = 18 + 42 - 60 = 0 so x = 3 is a solution

OpenStudy (anonymous):

so both x = 3 AND x= -10 would be solutions?

OpenStudy (campbell_st):

thats correct.... you can substitute x = -10 and check

OpenStudy (anonymous):

Ah, I see what you mean. Thanks a lot man!